### v1

# CollabStory: Multi-LLM Collaborative Story Generation and Authorship Analysis

Abstract
The rise of unifying frameworks that enable seamless interoperability of Large Language Models (LLMs) has made LLM-LLM collaboration for open-ended tasks a possibility. Despite this, there have not been efforts to explore such collaborative writing. We take the next step beyond human-LLM collaboration to explore this multi-LLM scenario by generating the first exclusively LLM-generated collaborative stories dataset called CollabStory. We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author stories. We generate over 32k stories using open-source instruction-tuned LLMs. Further, we take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for human-human multi-author writing tasks and analysis. We extend their authorship-related tasks for multi-LLM settings and present baselines for LLM-LLM collaboration. We find that current baselines are not able to handle this emerging scenario. Thus, CollabStory is a resource that could help propel an understanding as well as the development of techniques to discern the use of multiple LLMs. This is crucial to study in the context of writing tasks since LLM-LLM collaboration could potentially overwhelm ongoing challenges related to plagiarism detection, credit assignment, maintaining academic integrity in educational settings, and addressing copyright infringement concerns. We make our dataset and code available at https://github.com/saranya-venkatraman/multi_llm_story_writing.
CollabStory:
Multi-LLM Collaborative Story Generation and Authorship Analysis
Saranya Venkatraman, Nafis Irtiza Tripto, Dongwon Lee The Pennsylvania State University, USA {saranyav, nit5154, dongwon}@psu.edu
1 Introduction
Figure 1: CollabStory contains over $32k$ creative stories written collaboratively by up to 5 LLMs. Each story segment is generated by a single author, that then passes the narrative baton to the next, completing the storyline part by part in a sequential manner.
Generative Large Language Models (LLMs) are being used more widely and becoming ubiquitous in real-world scenarios. There is particular interest in understanding the use of such LLMs in various writing tasks as writing assistants or collaborators in machine-in-the-loop settings Yeh et al. (2024); Kim et al. (2023); Zhong et al. (2023); Singh et al. (2023); Yang et al. (2022); Yuan et al. (2022); Lee et al. (2022); Clark et al. (2018). So far though, this has only been explored in the case where a human is present. However given the rise of unifying frameworks that bring together and make LLMs from different sources interoperable, such as vLLM [^1] https://docs.vllm.ai/en/stable/, LangChain [^2] https://www.langchain.com/langchain, and HuggingFace [^3] https://huggingface.co/, the prospect of LLMs seamlessly collaborating and even handing off tasks to one another without external routing algorithms is on the horizon. This is particularly immediately possible with open-source models that are already being used by over 100K users per month (according to the number of downloads reported by HuggingFace). Despite the ease of interoperability of such LLMs, so far, automated writing assistants have been used only in collaboration with human authors or with a single LLM. Therefore, this study explores collaborative creative story-writing scenarios involving multiple LLMs, i.e. LLM-LLM collaboration.
Dataset # Stories # Authors  Avg words M-M Collaboration Available
STORIUM Akoury et al. (2020) 5,743 30,119 ~19k tokens H-H X
CoAuthor Lee et al. (2022) 830 58 418 H-M X
StoryWars Du and Chilton (2023) 40,135 9,494 367 H-H X X
CollabStory [Ours] 32,503 5 725 M-M
[…]
Collaborative creative story writing entails multiple authors contributing separate segments to form a coherent storyline (see Figure 1 for our dataset schema). Although individual LLMs excel at generating story plots, collaborative writing presents unique hurdles. Models must seamlessly continue the existing storylines generated so far by other models, even if they do not align perfectly with their own language distribution. The rise of multi-agent Artificial Intelligence (AI) underscores the potential for combining the expertise of agents specialized in various tasks. While previous mixture-of-experts scenarios focused on agents proficient in task-oriented settings Zhang et al. (2024); Pan et al. (2024); Liu et al. (2023); Li et al. (2023b, a), the emergence of LLMs conversing for continuous generative tasks in open domains is noteworthy. Imagine the possibilities when multiple LLMs collaborate; one LLM can generate compelling stories, but what if we put them together?
In this study, we attempt to address this question through a collaborative creative story-writing scenario involving multiple open-source LLMs. This is a crucial setting to study in the context of writing tasks since LLM-LLM collaboration could potentially overwhelm ongoing challenges related to plagiarism detection, credit assignment, maintaining academic integrity in educational settings, and addressing copyright infringement concerns.
We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author creative stories. This exploration is novel, as previous studies have primarily focused on human-LLM collaboration. Towards this goal, we generate the first multi-LLM collaborative story dataset called CollabStory using open-source LLMs. We select 5 frequently used LLMs (with number of downloads on HuggingFace for May 2024 provided in parenthesis): Meta’s Llama ($>540k$ downloads, Touvron et al. (2023)), Mistral.ai’s Mistral ($>1000k$ downloads, Jiang et al. (2023)), Google’s Gemma ($>180k$ downloads, Team et al. (2024)), AllenAI’s Olmo ($>26k$ downloads, Groeneveld et al. (2024)) and Microsoft’s Orca ($>22k$ downloads, Mukherjee et al. (2023)) to replicate a scenario in which commonly used LLMs from different organizations are being used in conjunction towards a single task. We demonstrate how one such dataset can be developed and the considerations involved in building an iterative Multi-LLM story-writer. We take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for multi-author writing tasks and analysis for human-human collaboration for over 15 years. We replicate their task settings and present baselines of different authorship-related tasks such as authorship verification and attribution for LLM-LLM collaboration and demonstrate that current baselines are challenged by this emerging scenario. CollabStory is the first resource that could help propel an understanding as well as the development of new techniques to discern the use of multiple LLMs in text.
Our work is motivated by the implications of Multi-LLM settings for different stakeholders (LLM developers, end-users) and considerations (such as credit assignment, legality of usage) arising in the generative AI landscape. As one example, a malicious actor might assemble texts from different LLMs together in one document to evade current detectors and successfully spread misinformation. Our discussion will further elaborate on the tasks our dataset enables and why it is crucial to develop methods to tackle the incoming challenges of machine-machine collaboration.
2 Related Work
LLMs as Collaborative Writers
LLMs are being increasingly used as writing assistants or to paraphrase, edit or enhance human-written written texts in machine-in-the-loop settings Kim et al. (2023); Singh et al. (2023); Yang et al. (2022); Clark et al. (2018). GhostWriter Yeh et al. (2024) and Wordcraft Yuan et al. (2022) are tools that enable users to co-write stories using instructions Yuan et al. (2022). Zhong et al. (2023) use “writing modes” as a control signal to better align the machine during co-writing with humans. CoAuthor positions GPT3.5 as a writing collaborator for over 50 human participants to co-write creative and argumentative stories Lee et al. (2022).
Datasets
Despite such emerging tools, only a handful have developed datasets that can be leveraged to understand collaborative story writing. One such resource is the STORIUM dataset Akoury et al. (2020) that contains over 5k creative stories written and obtained from human-human collaboration. In terms of human-machine co-writing, CoAuthor Lee et al. (2022) and CoPoet Chakrabarty et al. (2022) remain one of the few publically available datasets of human-machine collaborative creative story and poem writing, respectively. Beyond creative writing, Zeng et al. (2024) developed the first machine-human academic essay dataset as a means to study boundary detection for academic settings. A comparison of CollabStory with existing datasets is provided in Table 1.
# Number # Words per Author / # Author Order # Prompts per # Stories Authors HuggingFace distribution of LLMs used
of Authors ($N$) # Total Words Permutations Author Order
1 900 / 900 4 1800 7200 Gemma google/gemma-1.1-7b-it [^8] https://huggingface.co/google/gemma-1.1-7b-it
2 450 / 900 12 600 7200 Llama meta-llama/Llama-2-13b-chat-hf [^9] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf
3 300 / 900 15 480 7200 Mistral mistralai/Mistral-7B-Instruct-v0.2101010https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2
4 225 / 900 15 480 7200 Orca microsoft/Orca-2-13b [^11] https://huggingface.co/microsoft/Orca-2-13b
5 180 / 900 15 480 7200 Olmo allenai/OLMo-7B-Instruct [^12] https://huggingface.co/allenai/OLMo-7B-Instruct
Table 2: Summary of Data Collection Statistics
3 Methodology
3.1 CollabStory: Dataset Creation
We generate a dataset of creative stories using five open-source instruction-tuned LLMs: Llama2 Touvron et al. (2023), Olmo Groeneveld et al. (2024), Gemma Team et al. (2024), Mistral Jiang et al. (2023) and Orca Mukherjee et al. (2023) (model details are provided in Table 2). The main focus of our data generation is to simulate a scenario where LLMs from different sources (organizations) collaboratively work on a storyline, handing off control of the story from one LLM to the next. The stories in our dataset vary in the number of authors/LLMs involved, from being written entirely by a single LLM to written collaboratively by between 2 to up to all 5 LLMs. In this document, we refer to each of the LLMs as “authors". For cases where we refer to the human author, we specifically mention “human” author/writer. We generate our dataset by prompting various LLMs using creative writing prompts from an existing dataset called the Writing Prompts (WP) Dataset. The Writing Prompts Dataset was collected by Fan et al. (2018) using Reddit’s r/WritingPrompts/ forum that contains premises or prompts for stories. The WP dataset consists of a cleaned subset of story prompts and corresponding human-written stories using filtration criteria such as removal of stories that are bot-generated, less than 30 words long, contain profanity, general announcements, and so on. We used the test split [^4] https://www.kaggle.com/code/ratthachat/writingprompts-combine-one-line-data-for-gpt2/input?select=writingPrompts of this dataset as the source of prompts for LLM generated stories. We also filter out prompts that do not have at least one corresponding human-written story that is at least 800 words long. We do this to ensure that the prompt itself does not preclude longer storylines. We chose 800 words as a criteria as a means to include stories that are slightly longer than the average of the dataset. The average length (number of words) of articles in the test set is 675.75 words. Out of 15138 total prompt-story pairs, this left us with 4623 data points. For each prompt, we divide the total goal article length (800-900 words) by the number of authors ($N$) to calculate the length of each part or story chunk to be written by each author, such that the writing load is distributed roughly uniformly amongst the LLM authors. We also generate different permutations of LLM authorship order such that every author can contribute to random parts of the story and we ensure that our dataset does not have any spurious correlations between LLM/author and story sections such as the beginning, or ending. For each value of the numbers of authors i.e. $N\in\{1,2,3,4,5\}$, we generate all possible permutations of author orders. For example, for  N = 3, two examples of author order permutations could be:
$Olmo\rightarrow Mistral\rightarrow Llama$
$Gemma\rightarrow Llama\rightarrow Mistral$
From all such possible permutations, we sample the minimum of either total possible orders or 15 as the number of author orders. For each author order, we then generate stories using each of the prompts from a unique set of prompts per N. Our goal number of stories for each N was set to 7200 stories. A summary of the words written by each author, author order permutations, and prompts per author, as well as the pool of 5 authors and their corresponding model checkpoints used for generating all story parts is shown in Table 2.
3.2 LLM prompting
For each value of $N$, we used different prompts to generate story parts sequentially, as detailed in Table 3. Utilizing the vLLM library [^5] https://docs.vllm.ai/en/stable/, we accessed and generated text from various LLMs. Initially, we conducted a pilot study to refine our prompts by generating and reviewing 100 articles. For the "Beginning" prompt, the first LLM used only the original r/WritingPrompts/ input. For subsequent parts, we found that longer input prompts reduced story length, so we used Falcon.ai summarizer [^6] https://huggingface.co/Falconsai to condense the story so far into under 80 words, allowing LLMs to generate longer sequences. We also included the last sentence of the story so far for smooth continuity. Prompts for different sections only varied in their instructions to “begin”, “continue”, or “conclude” the story. Additionally, we added an instruction to prevent LLMs from generating extraneous instructions. More details are provided in Section A.2.
Story Part Prompt Template
Beginning You are a creative story writer. Write a story that starts with the prompt {starting prompt} in around {n} words. Do not add any instructions. Start the story as follows:
Middle Write {n} words to continue this storyline: {summary of story so far}. Continue from this sentence: {last sentence from previous part}
Ending Write {n} words to conclude this storyline: {summary of story so far}. Do not add any instructions. Continue from this sentence: {last sentence from previous part}
Table 3: Prompt templates for different parts of the story. {n} here denotes the number of target words for each author.
Part 1 Part 2 Part 3 Part 4 Part 5
Mistral Gemma Llama Olmo Orca
Years ago, in the depths of a dark and desolate forest, I made a promise to a witch. Desperate for her help to save my beloved firstborn son, I agreed to offer her my firstborn in return for her healing magic. The old crone’s cackling laughter still echoed in my ears as I made that fateful vow. But time passed, and my wife bore us three more children. With each birth, I was filled with relief and gratitude, knowing that our family had been spared the tragedy that had once threatened to tear us apart. Yet no matter how hard I tried, I couldn’t bring myself to fulfill my promise to the witch. Now, years have passed, and the witch has grown tired of waiting. Her messages, delivered by raven messenger, are becoming more and more urgent. The ground beneath my feet seems to shake with her anger, and I can’t help but fear for the consequences of my broken vow. Desperate for a solution, I turn to the only person who might be able to help me. But will she be able to save me from the wrath of the vengeful witch? Her pale, ethereal face held a cruel smile, her voice like the rustling of dry leaves in a desolate forest. "Your son’s life hangs in the balance, dear one," she hissed, her voice dripping with malice. "Yet, you have failed to fulfill your part of the bargain." Panic surged through me as I realized the witch was right. I had never presented her with a child. The burden of my desperation had become a heavier chains, tethering me to an unspoken debt. Desperation fueled the flames of my fear, as I knew that if I failed, not only would my son succumb to his fate, but the wrath of the vengeful witch would descend upon me. Time seemed to warp, stretching into an agonizing eternity as I considered my options. Could I find a way to appease her, to escape the consequences of my foolish decision? I paced back and forth in my chambers, my mind racing with thoughts of the witch and the debt I owed her. I had been so desperate to save my son, I had agreed to give her my firstborn child in exchange for her healing magic. Now, years had passed and the witch had grown tired of waiting. She had come to me, demanding that I fulfill my end of the bargain. I knew that I could not give her my son. He was all I had left of my beloved husband, and I would rather die than lose him. But I also knew that I could not defy the witch. She had the power to take my life, and the lives of my children, with just a wave of her hand. I spent hours pacing, trying to think of a way to appease her. But no matter what I did, I could not come up with a solution. I was trapped, and I knew it. The witch had me in her grasp, and she would not let me go. Finally, in desperation, I decided to seek out the help of a wise old woman who lived in the village. A few days ago, I came to her, my heart heavy with grief and my soul wounded. She, in turn, looked at me with a kind and gentle expression. Then, with a wave of her hand, she began her spell. I watched as the darkness in my heart began to fade, and my love for my children grew stronger than ever. She whispered words of power and love, and the bond between my wife and I was restored. From that day forward, we were a family once more, united by the magic of the wise old woman. The wise old woman had always been a renowned healer in our village. Her knowledge of herbs and potions was unmatched, and many came to her seeking relief from various ailments. I was one of those who sought her help, when my firstborn son fell gravely ill. His fever raged, and his tiny body was wracked with pain. Desperate for her help to save my beloved firstborn son, I agreed to give her my firstborn in exchange for her healing magic. I returned to my wife, heartbroken and ashamed, to tell her of the bargain I had made. But time passed, and my wife bore us three more children. With each birth, I was filled with relief and gratitude, knowing that our family had been spared the tragedy of losing a child. The wise old woman had honored her part of the bargain, and my son grew strong and healthy under her care. He became a skilled healer in his own right, following in the footsteps of the wise old woman who had saved his life.
Table 4: Example of a 5-part LLM story from CollabStory dataset for the prompt: “Years ago, you promised your firstborn to a witch. Since then, despite your best efforts, you can’t seem to give him away. The witch is starting to get pretty mad.
3.3 Post-processing and filtering
Though we used instruction-tuned LLMs, they do not follow instructions perfectly. Though our goal number of words per story was 800-900 words, we used the upper limit to calculate the number of words each LLM should generate. From our pilot study, we found that most LLMs were undershooting their target number of words in the instruction. We also filtered out all stories in which at least one part was under 50 words long. We also removed all extra spaces from the stories and any repetitions of the instructions in rare cases. We also filtered for some additional types of noises detailed in Section A.3. After this filtration, we were left with the following number of stories per $N\in[1,5]:{7164,7070,6093,6955,5221}$ for a total of $32,503$ stories. An example of one such story from our dataset can be read in detail in Table 4.
Feature N=1 (H) N=1 (M) N=2 N=3 N=4 N=5
# Words 1352.26 ± 425.11 725.03 ± 288.32 1090.67 ± 207.43 1154.44 ± 112.24 1091.99 ± 85.67 995.42 ± 74.20
# Sentences 84.23 ± 34.85 41.90 ± 61.94 60.49 ± 14.49 64.38 ± 11.57 59.29 ± 11.24 53.86 ± 9.41
# Words in sentence 17.82 ± 28.18 18.27 ± 3.92 18.42 ± 2.85 18.30 ± 2.45 18.81 ± 2.47 18.88 ± 2.51
Vocabulary richness 0.34 ± 0.05 0.40 ± 0.09 0.36 ± $0.05^{*}$ 0.36 ± $0.03^{*}$ 0.37 ± $0.03^{*}$ 0.39 ± 0.03
% of stopwords 31.26 ± 4.76 37.00 ± 4.77 37.77 ± 3.11 37.39 ± 2.67 37.71 ± 2.39 37.41 ± 2.39
Readability 80.28 ± 33.26 75.16 ± 9.83 75.34 ± 8.13 75.25 ± 7.80 74.37 ± 8.00 74.80 ± 8.14
Entropy 38.49 ± 12.25 26.03 ± 13.10 38.42 ± 7.93 40.30 ± 4.75 37.97 ± 4.07 34.69 ± 3.58
Coherence 0.38 ± 0.05 0.46 ± 0.06 0.46 ± 0.06 0.46 ± 0.04 0.46 ± 0.04 0.46 ± 0.04
Table 5: Comparison of descriptive features for articles with number of authors $N\in[1,5]$, as compared with human-written single-author stories for the same prompts. Here, “H” $\to$ “Human”; “M”$\to$ Machine. * represents statistical significance ($\textit{p}<0.01$) compared to human-written stories.
3.4 Descriptive Statistics
We report the average and standard deviation of number of words, sentences, vocabulary richness, readability scores, and coherence scores using TextDescriptives Library [^7] https://github.com/HLasse/TextDescriptives for all parts of the story, as well as for different numbers of authors ($N$) in Table 5. There is a statistically significant difference between LLM-coauthored and human-written stories in vocabulary richness measured by type-token-ratio (TTR) for $N\in[2,3,4]$. For all other measures, our dataset’s machine-generated stories follow similar distributions to human-written stories. We thus ensure the quality of our dataset using automated measures of readability and coherence by using human-generated stories as the reference text. Detailed statistics for each LLM are provided in the Appendix (Table 12).
4 Authorship Analysis: Extending PAN tasks for multi-LLM scenario
Plagiarism Analysis, Authorship Identification, and Near-Duplicate Detection, known as PAN tasks Bevendorff et al. (2023), have presented a persistent challenge, establishing benchmarks for analyzing multi-authored text among humans for more than 15 years. We extend the most common and repeated authorship-related tasks from the PAN multi-human-author task suite to the multi-LLM scenario. We then fine-tune and report performance using the following 5 baseline methods: Multinomial Naive Bayes (MNB) Losada and Azzopardi (2008), Support Vector Machine (SVM) Vapnik (1998), BERT Kenton and Toutanova (2019), ALBERT Lan et al. (2019), and RoBERTa Liu et al. (2019).
4.1 Task 1: Is a story written by multiple authors or not?
We randomly sample articles from the single-LLM authored stories i.e. $N=1$ as the negative class v/s articles from the multi-authored settings where $N\in[2,3,4,5]$ as the positive class. We sample from the single-LLM stories to keep the class distribution equal, based on the number of articles for each $N$. From Table 6, we see that for all methods, the performance at $N=5$ is higher than for $N=2$, gradually increasing with the value of $N$. Stories that have a higher number of authors are more distinct from single-authored ones. We conjecture that introducing more authors in the article might lead to more variations in the text, making stories with $N=5$ authors most easily distinguishable from stories without any such variations i.e. $N=1$.
Task 1: Multi-author or not
Method N=2 N=3 N=4 N=5
MNB 0.83 0.86 0.86 0.87
SVM 0.78 0.8 0.83 0.82
BERT 0.85 0.85 0.83 0.87
ALBERT 0.83 0.89 0.9 0.92
RoBERTa 0.88 0.92 0.9 0.93
AVG 0.83 0.86 0.86 0.88
Table 6: Performance is shown as F1-scores. AVG denotes average F1-score for each $N$. For each $N\geq 2$ we evaluate the classifiers on their ability to distinguish the stories from those written with $N=1$. Best performing method is in bold and second highest underlined for each N.
4.2 Task 2: How many authors have written a story?
The second task is to predict the number of authors involved in generating a story. For the CollabStory dataset, this means that class labels $\in[1,5]$. From Table 7 we see that the task of predicting exactly how many authors have co-written a story is easiest for $N=1$ in conjunction with findings from Task 1 that showed that multi-authored text can be more easily distinguished from single-authored text. Thus, here too it seems to be easiest to separate the single-authored texts from $N\geq 2$. However, for multi-authored stories, only BERT and RoBERTa perform better than other baselines (>0.72 F1), especially for $N\in[4,5]$. Overall, the performance across this task is low.
Task 2: Predict Number of Authors (N)
Method N=1 N=2 N=3 N=4 N=5
MNB 0.72 0.50 0.49 0.48 0.51
SVM 0.68 0.43 0.52 0.40 0.54
BERT 0.79 0.70 0.64 0.75 0.81
ALBERT 0.70 0.57 0.55 0.53 0.65
RoBERTa 0.76 0.68 0.63 0.72 0.74
AVG 0.73 0.58 0.57 0.58 0.65
Table 7: All scores are F1-scores. AVG denotes average F1-score for each $N$. This task is particularly challenging with only the single-authored stories ($N=1$) being correctly classified. For all multi-authored texts, BERT and RoBERTa perform better than others. Best performing method is in bold and second highest underlined for each N.
4.3 Task 3: Authorship Verification
This is a pair-wise sentence classification task where the goal is to predict if two adjacent sentences are written by the same author or not. For this task, we used all the sentences at LLM-LLM boundaries, that is the last sentence of part $i$ and the first sentence of part $i+1$. The negative class data samples were sampled as random pairs of consecutive sentences within each story part.
Task 3: Authorship Verification
Method N=2 N=3 N=4 N=5
MNB 0.65 0.64 0.60 0.63
SVM 0.63 0.62 0.62 0.63
BERT 0.73 0.73 0.73 0.71
ALBERT 0.92 0.89 0.89 0.89
RoBERTa 0.91 0.91 0.90 0.89
AVG 0.76 0.75 0.74 0.75
Table 8: All scores are F1-scores and AVG denotes average F-1 scores for each $N$ for the task of detecting authorship boundaries between sentence pairs. We see that generally, performance is slightly higher for $N=2$ across all classifiers. Best performing method is in bold and second highest underlined for each N.
From Table 8, we see that transformers-based fine-tuned methods perform well at this task. We also note that detecting sentence authorship boundaries seems to be slightly easier for the 2-author case than for $N\geq 3$.
4.4 Task 4: Authorship Attribution
Authorship Attribution involves predicting exactly who the author of a text article is. In the case of multi-LLM text, we design this task such that each data sample is homogeneous or each part is written by a single author and the classifier’s task is to identify its author.
Task 4: Authorship Attribution
N=1
Method Orca Olmo Llama Mistral Gemma AVG
MNB - 0.70 0.71 0.64 0.99 0.76
SVM - 0.61 0.68 0.58 0.97 0.71
BERT - 0.70 0.71 0.64 0.99 0.76
ALBERT - 0.78 0.73 0.70 0.99 0.80
RoBERTa - 0.73 0.70 0.68 0.99 0.78
N=2
Method Orca Olmo Llama Mistral Gemma AVG
MNB 0.49 0.51 0.52 0.51 0.92 0.62
SVM 0.51 0.55 0.54 0.59 0.79 0.62
BERT 0.54 0.54 0.63 0.58 0.95 0.68
ALBERT 0.56 0.58 0.63 0.59 0.96 0.69
RoBERTa 0.49 0.62 0.60 0.56 0.94 0.68
N=3
Method Orca Olmo Lama Mistral Gemma AVG
MNB - 0.60 0.67 0.63 0.94 0.71
SVM - 0.57 0.65 0.57 0.82 0.65
BERT - 0.58 0.69 0.67 0.95 0.72
ALBERT - 0.64 0.71 0.68 0.95 0.75
RoBERTa - 0.71 0.71 0.67 0.96 0.76
N=4
Method Orca Olmo Llama Mistral Gemma AVG
MNB - 0.58 0.65 0.63 0.91 0.69
SVM - 0.58 0.67 0.59 0.80 0.66
BERT - 0.59 0.70 0.68 0.93 0.73
ALBERT - 0.66 0.73 0.70 0.93 0.75
RoBERTa - 0.66 0.68 0.64 0.93 0.73
N=5
Method Orca Olmo Llama Mistral Gemma AVG
MNB 0.54 0.54 0.56 0.53 0.86 0.61
SVM 0.56 0.61 0.60 0.54 0.79 0.62
BERT 0.60 0.57 0.62 0.54 0.93 0.65
ALBERT 0.58 0.55 0.65 0.62 0.92 0.66
RoBERTa 0.56 0.69 0.61 0.58 0.92 0.67
Table 9: F1-scores for identifying the author of story parts across articles written by different numbers of authors. The 5 columns show each of the labels or authors. AVG denotes average F1-scores across all authors. Best performing method is in bold and second highest underlined.
From Table 9, we see that most of the authors seem hard to identify irrespective of the value of $N$, except for Gemma. We were expecting attribution to be easier the fewer the number of authors since the length of the parts contributed by each author would be longer. But there does not seem to be any such correlation in our dataset i.e. length contributed by each author does not correlate with their detection. To further explore why Gemma was easily identifiable, we manually inspected a random sample of 100 articles for discernible features or peculiarities in story parts written by Gemma v/s all other authors. We provide examples of story parts generated by Gemma and other LLMs in the Appendix (Table 11). We found that the text generated by Gemma is not noticeably different from that of other authors. We leave a deeper analysis of potential factors, such as training data-induced biases, story part, or author order for future work.
5 Discussion
PAN Task equivalent Task Description: Multi-LLM scneario Real-world Implications
Predict multi-author or not To determine if a text includes content from multiple LLMs or not Credit Assignment and Intellectual Property (IP) regulation
Predict number of authors To predict the number of LLMs involved in writing an article Keeping track of LLM-LLM agent interactions in growing open-source market
Author Verification To detect when authorship switches between LLMs To detect perjury, misinformation injection, falsifying editing in news articles, and text obfuscation
Authorship Attribution Predicting who wrote each text segment? Plagiarism detection
Style Change Detection and Attribution Finding all positions in the text where authorship changes and who wrote each segment Classroom settings: Academic Integrity, detecting use of multiple open-source and free-to-use LLMs to surpass detection methods
Table 10: Real-world implications of the tasks involved in understanding LLM-LLM collaboration for writing tasks
Recent developments have significantly advanced LLM-assisted writing, sparking widespread discussions about the nature of authorship. Beyond using LLMs for paraphrasing, editing, and enhancing text, there exists an extreme scenario where text is generated entirely by multiple LLMs. Our work addresses this extreme case, raising several nuanced authorship concerns: Who should be considered the true creative source in such a situation? Should all LLMs involved be credited? Or should the human developers designing the prompts be acknowledged as the primary authors? Moreover, should the LLM that contributed the most—whether in terms of word count, narrative depth, or plot twists—be granted greater ownership? There have been recent works addressing this question of ownership. For example, Joshi and Vogel (2024) and Dhillon et al. (2024) found that humans felt a sense of higher ownership when they wrote longer prompts and when the AI assistant-generated text length was shorter, respectively. The question of authorship when text is repeatedly paraphrased using LLMs has also been deeply considered by Tripto et al. (2023), who find that the notion of authorship is task-dependent and cannot be generalized.
These questions have profound implications for various stakeholders in the burgeoning socio-technical system of generative AI. Our research introduces authorship-related tasks using CollabStory, which can help address these concerns by accurately discerning the usage of multiple LLMs in texts. Our extension of PAN-inspired authorship tasks is closely linked to real-world implications, as follows:
Task 1: Predict multi-author or not
In the rapidly expanding and fiercely competitive market for LLMs, the ownership of content and the ability to prove the origins of creative work are becoming increasingly crucial. As the market evolves, closed-source LLMs are implementing stricter regulations and demanding credit assignment under various distribution licensing norms. In this context, the capacity to demonstrate that a text incorporates generated output from multiple LLMs is essential. This capability can effectively prevent any single stakeholder or developer from erroneously claiming exclusive rights to the content, thereby bolstering the defense against wrongful intellectual property (IP) claims.
Task 2: Predict number of authors
Predicting the exact number of LLMs involved in the writing process can help keep track of the frequency and extent to which LLMs are used collaboratively, as more and more models enter the open-source market. This is essential to understand whether such usage improves task performance or introduces inefficiencies beyond a certain threshold. Understanding the optimal number of LLMs or the degree to which LLMs can leverage each other’s strengths in writing tasks is vital. It ensures effective collaboration without unnecessary complexity, maximizing the benefits of combined model capabilities while avoiding overkill and collaboration for its own sake.
Task 3: Author Verification
With LLMs increasingly paraphrasing and editing each other’s texts, it becomes crucial to identify which spans were generated by different LLMs. Consider a scenario where a news article is paraphrased by one LLM and subsequently edited by another, with the latter introducing fallacies or misinformation. In such cases, discerning the contributions of each LLM is essential for identifying malicious LLM agents or the infiltration of critical content, such as media and news articles. This capability has significant applications, including detecting perjury and combating the adversarial obfuscation of text, thus maintaining the integrity and reliability of information.
Task 4: Authorship Attribution
Identifying the exact LLMs responsible for authoring a text is crucial for detecting and addressing plagiarism. This is particularly important in academic settings, where students might use closed-source LLMs without complying with content ownership and usage declaration regulations. This is possible also in cases where content from one LLM is being posed as that from another to claim higher ability or quality. An example of such a situation might be in a bid to motivate financial investors hoping to monetize and utilize LLMs for specific domains (such as medical applications, educational tools, and creative content generation).
6 Conclusion
In this work, we present CollabStory, the first exclusively LLM-LLM or machine-machine collaborative story dataset, and demonstrate the tasks it enables. We elaborate on why multi-LLM authorship tasks are crucial to study by discussing their real-world implications (summarized in Table 10). We will soon need “Catch As Catch Can” methods to not only find all points where authorship changes within an article (style change detection) but also simultaneously attribute each independent segment to the specific LLM author (attribution). As more and more LLMs are becoming easier to access, malicious actors could combine texts from different LLMs to evade automated and in-built misinformation flaggers, or students might circumvent credibility checks by having different LLMs write different sections of an academic article. Thus, CollabStory was developed as a resource with long-form stories written by multiple LLMs to support the development and expansion of tasks and methods that can help address incoming challenges brought by LLM-LLM interactions.
Limitations
Our work demonstrates one way of collecting a collaborative multi-LLM dataset. However, several variants are possible. Of course, as the LLM space is ever evolving, newer LLMs (e.g. Llama 3) became available as we were already collecting this dataset. Another aspect is that our dataset was collected in a uniform manner such that all LLMs contributed somewhat equal portions of text to a story. The next step would be to train a routing algorithm or a randomizer that could generate non-uniform collaborative texts. Our current analysis is unable to account for this setting and we leave this for future work. Additionally, the iterative generation process is resource-intensive and not easily scalable. We also acknowledge that LLM tasks beyond story writing are essential for a deeper understanding of how LLMs collaborate in open-ended generation tasks.
Ethics Statement
Using LLMs for creative story writing could relay some of the biases and harmful stereotypes present in the LLMs original training data since all our LLMs are trained on data from the internet. This is an important consideration before or during the dissemination of any such generated texts or stories. Transparency of the source of generated articles is important to avoid deception or wrongful content attribution. With creative writing tasks, it is also important to address any impact on creative professionals and guidelines to ensure that LLMs help enhance rather than undermine human creativity. We study LLM story-writing as a means to better prepare for a future of LLM-generated creative texts that might be misused in classroom settings, to manipulate public opinion on social media forums, and also to protect human writers against plagiarism amongst many other potential non-ethical usages.
Appendix A Appendix
A.1 Examples of story parts written by Gemma followed by other authors
We inspected 100 randomly sampled stories and inspected the parts written by Gemma in search of visibly discernible features or peculiarities that might have explained the ease of its detection as compared to other LLMs. To the naked eye, this text seemed to be similar to all the other parts of the story since we subjected all the story parts to the same filtration process. We suspect that other factors might be in play here such as the author order or tone differences, and leave this exploration to future attempts (Table 11 with examples of story samples in next page).
A.2 LLM Prompting
For for all $N\geq 2$, we provided the summary of the story so far as an input in the prompt. To make sure that our story parts had smooth continuity, we also used the last sentences of the story so far as input. This made sure that the generating LLM has access to the last sentence in addition to the overall storyline to continue the story as seamlessly as possible. This second input is denoted as “last sentence from previous part” in Table 3. Other than this, our prompt for the three types of story sections only differed in the instruction of writing either the “beginning”, “continue”, or “conclude” the storyline so far. We also had to add an instruction to stop the LLMs from generating any additional instructions as from our pilot study, we found that some LLMs (Orca and Llama) would often first generate a rephrasing or more detailed version of our instruction before generating the actual story content.
A.3 Dataset Cleaning
For each prompt, we gave each LLM 20 maximum attempts to re-generate that particular story part if it fell 15 or more words shorter than the goal length in the previous iteration. Despite this, we had instances of very short story parts that would have made the average article length too short or led to a very skewed representation of one LLM v/s the rest. Thus, we discarded such stories. Additionally, we were able to notice two formatting peculiarities for Gemma and LLama. Particularly, Gemma’s story parts often began with a short title for the section it was to generate surrounded by “###” for example “### The return of the Jedi ###”. Llama on the other hand was appending a “The end” whenever it was its turn to write the ending part of a story. We removed all cases of these two substrings using regular expressions search and deletion as a means to unify the flow of the story across all LLMs and to make sure particular LLMs weren’t identifiable only due to such formatting details. We also removed all extra spaces from the stories and any repetitions of the instructions in rare cases.
Gemma Not Gemma
The chase has begun, and adrenaline floods his veins. The blinding headlights pierce through the night, illuminating the frantic dance of tires on asphalt. The adrenaline-fueled chase weaves through the labyrinth of city streets, sirens screaming like tormented spirits. The driver, hands white on the wheel, desperately maneuvers, adrenaline masking the fear gnawing at his conscience. Every corner, every alleyway, feels like a potential trap, each sudden movement a potential assault. The police officer, unwavering in his pursuit, his seasoned instincts guiding him through the chaos. He maneuvers his cruiser with precision, anticipating the driver’s next move. His heart pounds in his chest, adrenaline battling with the growing sense of urgency. He knows the city like the back of his hand, weaving through traffic, his eyes glued to the fleeing vehicle. As the chase intensifies, the driver’s desperation escalates. He risks collisions, weaving through pedestrians, pushing the limits of his vehicle’s performance. The officer remains calm, focused, his training taking over. He knows he must bring this chase to an end, but he also understands the risks involved. Swallowing his fear, he decides to take the next exit and lead his pursuers through the winding, narrow streets of the old town. The streets are alive with the sound of market vendors calling out their wares, children laughing, and the faint sounds of a distant band playing music. The narrow, cobblestone alleys are lined with ancient buildings, each one more beautiful than the last. He knows that this labyrinth of streets will give him an advantage, but it also increases the risk of colliding with a pedestrian or another car. Despite the danger, he continues to navigate the maze of alleys, turning sharply at each intersection. He can hear the revving of engines and the screeching of tires behind him, but he refuses to look back. He knows that every twist and turn could be his last, but he is determined to outsmart his pursuers and escape. As he rounds the final corner, he sees the safety of the forest beyond the old town walls. He grits his teeth, puts the pedal to the metal, and bursts through the gates, leaving the chaos of the chase behind him.
Emily stood frozen, her heart pounding in her chest. A sense of awe washed over her as she approached the base of the tree and noticed a faint glow emanating from beneath the ground. Curiosity overwhelmed her, and she cautiously lowered a hand into the soft earth. As her fingers brushed against something cold and smooth, a surge of energy flowed through her body. A deep voice rumbled in her mind, "You have awakened. The spirit of the wolf within you has been awakened." The earth trembled beneath her feet, and a swirling mist began to envelop Emily. She felt her bones shifting, her muscles growing taut. Her reflection in the nearby pond transformed from a girl to a magnificent silver wolf. She barked, a primal call echoing through the woods. The glow from beneath the oak tree intensified, and a majestic wolf spirit emerged, its eyes glowing with wisdom and power. The spirit guided Emily through the forest, teaching her the secrets of her newfound lineage. She learned of the ancient bond between wolves and humans, the importance of loyalty and bravery. As she embraced her new identity, she found herself a part of an intricate web of relationships. Her new pack, made up of both humans and wolves, accepted her with open arms. Together, they faced challenges, both natural and supernatural. Emily’s courage and quick thinking saved the pack on more than one occasion. Her transformation brought her closer to her true self, and she no longer felt lost or alone. The bonds of loyalty and friendship grew stronger as they worked together to protect their territory. Emily felt a deep sense of belonging, and she knew that she had found her place in the world. The woods, once a place of solitude, had become her home, a place of love and acceptance. Emily’s life as a wolf shifter was not without its difficulties, but she faced each challenge with the same bravery and determination that had always defined her. She had found her true identity, and nothing would ever shake her sense of belonging or her commitment to her pack. As the sun set over the glade, Emily looked up at the stars, her heart filled with gratitude and love.
In the unspoken words hanging between us, Sarah and I understand that the journey to healing will be arduous. There will be moments of doubt, days when the weight feels unbearable, when the waves of sadness threaten to pull us under. But we are not in this together. We are stronger than our individual struggles. Together, we will learn to navigate the storms, to embrace the sunshine, to rediscover who we are beyond the baggage we once carried. Sarah’s smile brightens my day, and I know that this new chapter in our lives is going to be filled with adventure, growth, and most importantly, love. We’ll stumble, we’ll trip, but we’ll always have each other to pick us back up and move forward, stronger than before. I take Sarah’s hand and squeeze it tightly.
Her words seemed to hang in the air like acrid smoke, but I knew better. The voice, the chilling certainty in its delivery, had belonged to something far more sinister. I glanced at the others, their faces pale, their eyes wide with disbelief. "There’s no way this is real," one whispered, her voice trembling. But the undeniable presence of death in our living room told us otherwise. The figure shifted, its skeletal hands brushing the floor with an unholy whisper. "Your time has come," it hissed, its voice like nails on a chalkboard. Panic surged through me, my limbs frozen in terror. The others stumbled around, scrambling for escape, but their movements were clumsy, their screams muted by fear. In that moment, I realized that the only way out was through the front door, directly into the path of the creature that had come for me. With a surge of adrenaline, I broke free from the group, sprinting towards the exit. As I reached the door, it swung open to reveal Tia, the figure from my nightmares, standing there. But instead of fear, I felt relief. She had a warm smile on her face, and her eyes were full of compassion. "You have been chosen to bear the mark of the guardian," she explained, holding out the hourglass. "I have protected this town for generations, and now it’s your turn to carry on the tradition." I hesitated for a moment, but then I accepted the hourglass with a sense of pride and purpose. Together, Tia and I faced the darkness that lurked outside, ready to protect our town and its people from the evil that sought to harm them. From that day on, I became the guardian of my town, using the powers granted to me by Tia to keep the peace and protect those around me. The people of the town were grateful, and I knew that I had found a new purpose in life.
Table 11: As can be seen from a few examples, the text written by Gemma is not necessarily different from that of other authors (in the second column). Thus, from our manual inspection, we were unable to attribute any noticeable features to the higher authorship attribution performance for Gemma.
Feature Author K=1 K=2 K=3 K=4 K=5
# Words Gemma 172.97 ± 16.47 157.17 ± 36.22 124.51 ± 46.13 129.76 ± 47.16 133.75 ± 46.90
Llama 172.51 ± 19.91 170.88 ± 13.28 173.35 ± 15.29 174.76 ± 16.18 172.23 ± 19.33
Mistral 177.25 ± 12.61 182.24 ± 12.88 178.09 ± 22.71 178.82 ± 19.47 178.15 ± 22.69
Olmo 168.01 ± 8.69 197.91 ± 18.93 194.89 ± 23.64 192.64 ± 26.60 191.41 ± 30.28
Orca 174.45 ± 22.20 175.61 ± 10.42 178.11 ± 15.62 178.01 ± 14.37 177.89 ± 16.57
Lexical Diversity Gemma 0.67 ± 0.04 0.67 ± 0.05 0.70 ± 0.07 0.69 ± 0.07 0.68 ± 0.06
Llama 0.60 ± 0.05 0.61 ± 0.05 0.59 ± 0.05 0.58 ± 0.05 0.58 ± 0.06
mistral 0.64 ± 0.04 0.62 ± 0.04 0.61 ± 0.05 0.61 ± 0.04 0.61 ± 0.05
Olmo 0.62 ± 0.05 0.57 ± 0.06 0.58 ± 0.06 0.58 ± 0.06 0.57 ± 0.07
Orca 0.62 ± 0.05 0.62 ± 0.04 0.61 ± 0.04 0.61 ± 0.05 0.60 ± 0.05
Readability Gemma 75.95 ± 8.53 77.13 ± 9.70 75.13 ± 12.29 74.11 ± 11.81 73.28 ± 12.96
Llama 83.11 ± 8.35 82.61 ± 8.50 82.75 ± 9.00 80.88 ± 10.10 80.13 ± 9.19
Mistral 81.04 ± 8.58 83.99 ± 8.40 81.59 ± 10.07 82.02 ± 9.09 79.91 ± 9.44
Olmo 80.78 ± 9.01 83.31 ± 9.87 81.41 ± 9.81 80.55 ± 10.73 80.45 ± 10.20
Orca 83.08 ± 8.54 82.51 ± 8.45 80.86 ± 9.65 79.95 ± 9.97 79.51 ± 9.80
Coherence Gemma 0.49 ± 0.07 0.47 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.48 ± 0.08
Llama 0.44 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.49 ± 0.08
mistral 0.46 ± 0.07 0.44 ± 0.07 0.45 ± 0.07 0.45 ± 0.07 0.46 ± 0.07
Olmo 0.47 ± 0.08 0.43 ± 0.07 0.45 ± 0.08 0.45 ± 0.08 0.45 ± 0.08
Orca 0.44 ± 0.07 0.45 ± 0.08 0.45 ± 0.08 0.46 ± 0.08 0.47 ± 0.07
Table 12: Descriptive Statistics or Features for stories generated by different authors for different parts of the stories. Here, “K” represents the part of the story written, i.e. K=1 corresponds to the first part of the story, K=2 referees to the second part, and so on.